# User Feedback Alignment for LLM-powered Exploration in Large-scale Recommendation Systems

### v3

@@ Abstract
Exploration, the act of broadening user experiences beyond their established preferences, is challenging in large-scale recommendation systems due to feedback loops and limited signals on user exploration patterns. Large Language Models (LLMs) offer potential by leveraging their world knowledge to recommend novel content outside these loops. A key challenge is aligning LLMs with user preferences while preserving their knowledge and reasoning. While using LLMs to plan for the next novel user interest, this paper introduces a novel approach combining hierarchical planning with LLM inference-time scaling to improve recommendation relevancy without compromising novelty. We decouple novelty and user-alignment, training separate LLMs for each objective. We then scale up the novelty-focused LLM’s inference and select the best-of-n predictions using the user-aligned LLM. Live experiments demonstrate efficacy, showing significant gains in both user satisfaction (measured by watch activity and active user counts) and exploration diversity.
@@
Jianling Wang1, Yifan Liu2, Yinghao Sun3, Xuejian Ma2, Yueqi Wang2, He Ma2, Steven Su2, Ed H. Chi1, Minmin Chen1, Lichan Hong1, Ningren Han2, Haokai Lu1 1Google DeepMind  2YouTube  3Google Labs {jianlingw, yifanliu, sunmo, xuejianma, yueqiw, htm, susteven, edchi, minminc, lichan, peterhan, haokai}@google.com
@@ 1 Introduction
Large Language Models (LLMs) present a significant opportunity to revolutionize recommendation systems Wu et al. (2024), due to their powerful reasoning, planning, and world knowledge capabilities. Traditional recommendation backbones, such as collaborative filtering and content-based methods, typically suggest items by identifying similar users based on past interactions, which often reinforce existing preferences and perpetuate feedback loops Chaney et al. (2018); Mansoury et al. (2020). LLMs can overcome these limitations by leveraging their vast world knowledge to generate recommendations that go beyond a user’s historical interactions, introducing novel and diverse items, to drive long-term user engagement Chen (2021) and reduce reliance on past behavior alone.
@@
Among recent advancements leveraging LLMs for recommendation systems Bao et al. (2023); Lin et al. (2024a); Wang et al. (2024a); Christakopoulou et al. (2023), the hierarchical planning paradigm Wang et al. (2024b) stands out a promising and deployable approach that combines LLMs for high-level guidance with traditional recommenders for efficient item-level serving. As this solution has been adopted in industry, the next challenge is integrating real-world human feedback into the LLM. While human feedback is key to optimizing LLMs Ouyang et al. (2022), systematically incorporating it into recommendation systems remains an under-explored area, offering both challenges and opportunities for future research.
@@
Effectively using real-world human feedback is challenging because recommendation systems rely on noisy implicit signals (e.g., clicks or dwell time) instead of explicit comparative judgments (e.g., side-by-side comparisons). This make it hard to translate such feedback into robust training objectives for LLMs that align with users’ true preferences. More importantly, balancing novelty and relevance, two usually competing objectives, is crucial for exploration in recommendation systems as relevant novel contents drive sustained user satisfaction. Initial experiments with the hierarchical planning Wang et al. (2024b) framework, using an LLM as a novelty model to identify novel interest cluster and subsequently retrieve relevant items, demonstrated the potential of this approach. However, aligning the novelty model’s predictions with user preferences remains challenging. Directly fine-tuning with more users’ interaction history data yielded neutral results, also raised concerns about memorization and loss of novelty. Attempts at RLHF Ouyang et al. (2022) with a reward model also proved unsuccessful as it undermine the controlled generation capability (see in Sec. 3).
@@
To address these challenges, we propose a novel, decomposed approach that leverages two specialized LLM models for high-level planning: a novelty model and an alignment model. To balance novelty and relevance, the alignment LLM is trained specifically to rate the novelty model’s predictions based on observed user feedback. This separation allows for the independent optimization of novelty generation and preference alignment. Moreover, to further improve the system’s ability to generate relevant novel predictions, we scale inference-time compute by generating multiple independent predictions from the novelty model using a high temperature setting. The alignment model then acts as a selector, choosing the most user-aligned outputs from the novelty model. This combination of specialized models, collective user behaviors as training signal, and repeated sampling significantly increases the likelihood of generating recommendations that are both novel and relevant.
@@
In summary, this paper presents the following key contributions: (1) Collective User Feedback Alignment: We introduce an LLM-based alignment model specifically trained to evaluate the novelty model’s predictions based on collective user behaviors. By aggregating implicit signals like clicks and dwell time across user clusters, we enable the system to learn user preferences with reduced noise and bias. (2) Inference-Time Scaling: We demonstrate the effectiveness of repeated sampling at inference time, allowing the alignment model to select the most relevant predictions from a diverse set of candidates generated by the novelty model, thereby improving exploration. (3) Decomposed Novelty and Preference Modeling: We propose a novel paradigm that decouples novelty generation and preference modeling into two specialized LLMs. This separation allows for independent optimization of each objective, resulting in a significantly improved operating curve for user interest exploration by directly addressing the core challenge of balancing novelty with relevance through specialized models. The system is deployed on a commercial short-form video recommendation platform serving billions of users.
@@ 2 Related Work
LLMs for Recommendation Systems. The advances in LLM capabilities recently has drawn a lot of attention to their potential in recommendation systems Bao et al. (2023); Dai et al. (2023); Geng et al. (2023); Hou et al. (2023); Li et al. (2023); Liu et al. (2023a). One promising direction involves augmenting traditional recommendation models with LLM-powered feature engineering, including supplementary textual features or embeddings encoded world knowledge Xi et al. (2024); Ren et al. (2024); Liu et al. (2023b). Another approach focuses on directly generating recommendations using LLMs. For instance, Hou et al. and Gao et al. have experimented with prompting off-the-shelf LLMs to produce ranked lists of recommendations. Meanwhile, there are also work involves fine-tuning LLMs Singh et al. (2024); Bao et al. (2023); Lin et al. (2024b) to better align them with the recommendation domain, whether through incorporating domain-specific knowledge, generating new tokens, or predicting user preferences for specific user-item pairs. However, few of these methods are truly equipped to handle query-per-second (QPS) requirements of real-time applications. Wang et al. (2024a) addresses this by employing LLMs as data augmentation tools for conventional recommendation systems during training, thereby boosting performance without incurring additional serving costs.
Recommendation Exploration. Improving user interest exploration is key to broadening preferences and fostering long-term engagement Chen et al. (2021); Chen (2021); Su et al. (2024). However, a key challenge lies in the inherent closed-loop nature of existing recommendation systems Chaney et al. (2018); Mansoury et al. (2020); Chen et al. (2021). Training data is primarily derived from past user-item interactions, limiting the system’s ability to explore truly novel interests. While methods like PIE Mahajan et al. (2023) offer improvements through user-creator affinity and online bandit formulations, they remain confined by the system’s internal knowledge Chen et al. (2021). Building on the LLM-powered hierarchical architecture od Wang et al. (2024b), which guides user interest exploration at the cluster level, we focus on enhancing its performance through user feedback alignment. Our work represents an early investigation of effective user feedback signals and their integration into LLMs for recommendation systems.
@@ 3 Preliminaries
Hierarchical Planning Paradigm. In the hybrid hierarchical planning paradigm Wang et al. (2024b), LLMs focus on high-level planning by predicting novel user interests at the interest cluster level. Interest clusters are topically coherent item clusters generated from item metadata and content embedding Chang et al. (2024). To grant LLM with domain knowledge of our system, we finetuned the LLM using the novel interest transition patterns mined from users’ interaction history.
@@
As illustrated in Figure 1, during the high-level planning, given a user’s recent interaction history, represented as a sequence of $K$ clusters $S_{u}$, with $S_{u}\subseteq\mathcal{C}$ and $|S_{u}|=k$, the LLM predicts the next novel cluster $C_{n}$ for this user. Because online serving the LLM for a billion-user system is prohibitively costly, we pre-compute and store potential next interest transitions for all combinations of sampled k clusters $\mathcal{S}=\{S\mid S\subseteq\{C_{1},C_{2},\dots,C_{N}\},|S|=k\}$. During online serving, a user’s history is mapped to the corresponding pre-computed novel interest through looking up the precomputed interest transitions. At the lower level, a conventional, transformer-based sequential recommender backbone handles the computationally intensive task of item-level selection. However, instead of searching the entire item space, the backbone is constrained to recommend items only within the novel interest cluster $C_{n}$ identified by the LLM. This constraint combines the personalization capabilities of the backbone with the novelty-seeking behavior of the LLM, leading to a personalized recommendation experience enriched with serendipitous discoveries.
@@
We’ve launched this user interest exploration paradigm to the production recommendation system, which resulted in a rare combination of high novel item ratio and user satisfaction gain. The lightweight finetuning (<8k training examples) played a key role in preserving the LLM’s own knowledge while teaching it some understanding of our users’ interaction patterns.
@@
Limitation. The lightweight finetuning, however, has limitations: 1) The 8k training examples represented a limited view of the behavior of our large user base. 2) For the cluster combinations that are hard to reason, LLM has low prediction confidence, indicated by the novel interest predictions that don’t have a logical connections to the users’ existing interests. This hurts the relevancy of the recommendation thus user satisfaction.
Figure 1: Hierarchical planning paradigm.
@@
To improve the relevancy of the novel interest prediction, we initially try to increase the number of training examples, but novelty metrics are not sensitive to the change in A/B testing. Furthermore, due to the LLM’s tendency to repeat training data (our analysis showed a 40% chance of repetition during inference), scaling to more training example mined from the user history risked reinforcing the system feedback loop, impairing LLM’s ability to make novel recommendations.
@@
To align the LLM with user preference without amplifying the system feedback loop, we leverage live-traffic users’ feedback to LLM’s own recommendations, such as clicks, dwell time and repeated interaction, which is independent of the system behavior. We first tried the classic RLHF setup: RL finetune the novelty LLM directly with a reward model trained with user preference. However, this always resulted in the model quickly collapsing: 1) loss of controlled generation: after 5k steps, the LLM’s chance of predicting in the correct format drops from 99+to 2%; 2) Reward hacking: the model learned the high reward words, e.g. cat, BTS, toys, etc, and frequently predicts those words. While RLHF is effective in free form text generation in conversation settings, it turns out for structured tasks with strict format and content vocab requirement, a reward model is insufficient and cannot capture the nuanced task requirements and guide the RL finetuning process accordingly.
@@ 4 Method
To address the challenges in the classic RLHF, we introduce an inference-time scaling method Brown et al. (2024) with a decoupled dual-specialization modeling approach. Instead of directly fine-tuning the policy model (i.e., the novelty model responsible for planning the next cluster) through SFT or RLHF, our approach first performs independent sampling from the novelty model. This generates a diverse set of candidate interest clusters. And subsequently the best-of-n clusters are selected using a separate alignment model trained on collective user feedback based on their likelihood to resonate with users.
@@ 4.1 Preference Alignment on User Feedback
Aggregating Collective Human Feedback. Through per-query logging inside our LLM-powered recommender serving live traffic (detailed in Section 3, ‘Novelty model’ hereafter), we collect users’ preferences on LLM’s predictions. Specifically, for each predicted cluster $C_{n}$, we log the cluster sequence {$C_{1}$, …, $C_{K}$} used to represent the user, and the user’s feedback on $C_{n}$ (e.g., positive playback, like, share, skip, etc) . We then aggregate the feedback for each ({$C_{1}$, …, $C_{K}$}, $C_{n}$) pair, resulting in user preference training example denoted as ({$C_{1}$, …, $C_{K}$}, $C_{n}$ , $L_{(1,k),n}$). Here, $L_{(1,...,k),n}$ represents the aggregated user feedback score(e.g. like rate, share rate) for this particular interest cluster transition – that is, serving interest cluster $C_{n}$ to a user with historical viewing pattern represented by {$C_{1}$, …, $C_{K}$}).
We then post-process the aggregated feedback to: 1) normalize the feedback score, which can be skewed to very small value because the feedbacks, e.g. like, share, etc, are sparse. 2) filter cluster transition pairs with few user feedback. 3) round the feedback score to a fixed interval to account for margin of error in the aggregated stats.
@@
Alignment Reward Model Training. To align with collective user feedback, we trained an alignment reward model (‘alignment model’ hereafter) which scores on $C_{n}$ , using a cross-entropy loss between the prediction and the user’s actual aggregated engagement metric (i.e., positive playback rate). This alignment model is an LLM with the last layer being a linear projection layer.
@@
In Figure 2, we showcase a sample prompt describing users with {photography, Visual Effects, Special effects} and {dogs} interest clusters(assuming $K=2$). Collectively, those users prefer label 1 {wildlife, nature documentary} over label 2 {boats} as expressed in the feedback scores. Given two new labels, the trained alignment model also effectively assigns high preference score to {cats, video game, internet meme} over less relevant next interest cluster. These intuitive examples demonstrate the feasibility and potential of the alignment training.

### v4

@@ Abstract
Exploration, the act of broadening user experiences beyond their established preferences, is challenging in large-scale recommendation systems due to feedback loops and limited signals on user exploration patterns. Large Language Models (LLMs) offer potential solutions by leveraging their world knowledge to recommend novel content outside these loops. A key challenge is aligning LLMs with user preferences while preserving their knowledge and reasoning. To enhance planning for new user interests using LLMs, this paper introduces a novel approach that combines hierarchical planning with LLM inference-time scaling. It aims to improve recommendation relevancy without compromising novelty. We decouple novelty and user-alignment, training separate LLMs for each objective. We then scale up the novelty-focused LLM’s inference and select the best-of-n predictions using the user-aligned LLM. Live experiments demonstrate efficacy, showing significant gains in both user satisfaction (measured by watch activity and active user counts) and exploration diversity.
@@
Jianling Wang1∗, Yifan Liu2∗, Yinghao Sun3, Xuejian Ma2, Yueqi Wang2, He Ma2, Zhengyang Su2, Minmin Chen1, Mingyan Gao2, Onkar Dalal2, Ed H. Chi1, Lichan Hong1, Ningren Han2, Haokai Lu1 1Google DeepMind  2YouTube  3Google Labs {jianlingw, yifanliu, sunmo, xuejianma, yueqiw, htm, susteven, minminc, mingyan, onkardalal, edchi, lichan, peterhan, haokai}@google.com
**footnotetext: indicates equal contribution
@@ 1 Introduction
Large Language Models (LLMs) present a significant opportunity to revolutionize recommendation systems Wu et al. (2024), due to their powerful reasoning, planning, and world knowledge capabilities. Traditional recommendation backbones, such as collaborative filtering and content-based methods, typically suggest items by identifying similar users based on past interactions, which often reinforce existing preferences and perpetuate feedback loops Chaney et al. (2018); Mansoury et al. (2020). LLMs can overcome these limitations by leveraging their vast world knowledge to generate novel and diverse recommendations that go beyond a user’s historical interactions, thus driving long-term user engagement Chen (2021).
@@
Among recent advancements leveraging LLMs for recommendation systems Bao et al. (2023); Lin et al. (2024); Wang et al. (2024a), the hierarchical planning paradigm Wang et al. (2024c) stands out as a promising and deployable approach that combines an LLM, which provides high-level guidance, with traditional recommenders for efficient item-level serving. As this solution has been adopted in industry, the subsequent challenge lies in effectively integrating real-world human feedback into the LLM. While human feedback is key to optimizing LLMs Ouyang et al. (2022), systematically incorporating it into recommendation systems remains an under-explored area, offering both challenges and opportunities for future research.
@@
Using real-world human feedback is challenging because recommendation systems rely on noisy implicit signals (e.g., clicks or dwell time) instead of explicit comparative judgments (e.g., side-by-side comparisons). This makes it hard to translate such feedback into robust training objectives for LLMs that align with users’ true preferences. More importantly, balancing novelty and relevance – two usually competing objectives – is crucial for exploration in recommendation systems as relevant novel content drives sustained user satisfaction. Initial experiments with the hierarchical planning Wang et al. (2024c) framework, using an LLM as a novelty model to identify novel interest clusters and subsequently retrieve relevant items, demonstrated the potential of this approach. However, aligning the novelty model’s predictions with user preferences remains challenging. Directly fine-tuning with more users’ interaction history data yielded neutral results and raised concerns about memorization and loss of novelty. Attempts at RLHF Ouyang et al. (2022) with a reward model also proved unsuccessful as it undermined the controlled generation capability (see in Sec. 3).
@@
To address these challenges, we propose a novel, decomposed approach that leverages two specialized LLMs for high-level planning: a novelty model and an alignment model. To balance novelty and relevance, the alignment LLM is trained specifically to evaluate and rate the predictions of the novelty model based on observed user feedback. This separation allows for the independent optimization of novelty generation and preference alignment. Moreover, to further improve the system’s ability to generate relevant novel predictions, we scale inference-time compute by generating multiple independent predictions from the novelty model using a high temperature setting. The alignment model then acts as a selector, choosing the most user-aligned outputs from the novelty model. This combination of specialized models, training signals derived from collective user behaviors, and repeated sampling significantly increases the likelihood of generating recommendations that are both novel and relevant.
@@
In summary, this paper presents a system that has been deployed on a commercial short-form video recommendation platform serving billions of users. The key contributions are: (1) Collective User Feedback Alignment: We introduce an LLM-based alignment model specifically trained to evaluate the novelty model’s predictions based on collective user behaviors. By aggregating implicit signals(e.g. clicks and dwell time) for interest clusters transition across many users, we enable the system to learn user preferences with reduced noise and bias. (2) Inference-Time Scaling: We demonstrate the effectiveness of repeated sampling at inference time, allowing the alignment model to select the most relevant predictions from a diverse set of candidates generated by the novelty model, thereby improving exploration. (3) Decomposed Novelty and Preference Modeling: We propose a novel paradigm that decouples novelty generation and preference modeling into two specialized LLMs. This separation enables independent optimization for each objective. Consequently, it directly addresses the core challenge of balancing novelty with relevance via specialized models, leading to a significantly improved operating curve for user interest exploration.
@@ 2 Related Work
This research builds upon two primary streams of existing work: the application of LLMs to recommendation systems and the ongoing efforts to improve recommendation exploration.
LLMs for Recommendation Systems. The advances in LLM capabilities have recently drawn a lot of attention to their potential in recommendation systems Bao et al. (2023); Geng et al. (2023); Hou et al. (2024); Li et al. (2023); Liu et al. (2023); Wang et al. (2024b). One promising direction involves augmenting traditional recommendation models with LLM-powered feature engineering, including supplementary textual features or embeddings that encode world knowledge Xi et al. (2024); Ren et al. (2024). Another approach focuses on directly generating recommendations using LLMs; e.g., Hou et al. and Gao et al. have experimented with prompting off-the-shelf LLMs to produce ranked lists of recommendations. Meanwhile, there are also work involving fine-tuning LLMs Singh et al. (2024); Bao et al. (2023); lin2024rella to better align them with the recommendation domain, whether through incorporating domain-specific knowledge, generating new tokens, or predicting user preferences for specific user-item pairs. However, few of these methods are truly equipped to handle query-per-second (QPS) requirements of real-time applications. Wang et al. (2024a) addresses this by employing LLMs as data augmentation tools for conventional recommendation systems during training, thereby boosting performance without incurring additional serving costs.
Recommendation Exploration. Improving user interest exploration is key to broadening preferences and fostering long-term engagement Chen et al. (2021); Chen (2021); Su et al. (2024). However, a key challenge lies in the inherent closed-loop nature of existing recommendation systems Chaney et al. (2018); Mansoury et al. (2020); Wang et al. (2023). Training data is primarily derived from past user-item interactions, limiting the system’s ability to explore truly novel interests. While methods like PIE Mahajan et al. (2023) offer improvements through user-creator affinity and online bandit formulations, they remain confined by the system’s internal knowledge Chen et al. (2021). Building on the LLM-powered hierarchical planning architecture Wang et al. (2024c), which guides user interest exploration at the cluster level, we focus on enhancing its performance through user feedback alignment. Our work investigates the integration of effective user feedback signals into LLMs for recommendation systems.
@@ 3 Preliminaries
Hierarchical Planning Paradigm. In the hybrid hierarchical planning paradigm Wang et al. (2024c), LLMs focus on high-level planning by predicting novel user interests at the interest cluster level. Interest clusters are topically coherent item clusters generated from item metadata and content embedding Chang et al. (2024). To provide the LLM with domain knowledge of our system, we fine-tuned the LLM using the novel interest transition patterns mined from users’ interaction history.
@@
As illustrated in Figure 1, during the high-level planning, given a user’s recent interaction history, represented as a sequence of $K$ clusters $S_{u}$ (i.e., $|S_{u}|=k$), the LLM predicts the next novel cluster $C_{n}$ for this user. Because online serving the LLM for a billion-user system is prohibitively costly, we pre-compute and store potential next interest transitions for all combinations of sampled k clusters $\mathcal{\textbf{S}}=\{S\mid S\subseteq\{C_{1},C_{2},\dots,C_{N}\},|S|=k\}$. During online serving, a user’s history is mapped to the corresponding pre-computed novel interest through looking up the precomputed interest transitions. At the lower level, a conventional, transformer-based sequential recommender backbone handles the computationally intensive task of item-level selection. However, instead of searching the entire item space, the backbone is constrained to recommend items only within the novel interest clusters $C_{n}$ identified by the LLM. This constraint combines the personalization capabilities of the backbone with the novelty-seeking behavior of the LLM, leading to a personalized recommendation experience enriched with serendipitous discoveries.
@@
We’ve launched this user interest exploration paradigm to the production recommendation system, which resulted in a rare combination of high novel item ratio and user satisfaction gain. The lightweight finetuning (<8k training examples) was key to preserving the LLM’s pre-trained knowledge while imparting an understanding of our users’ interaction patterns.
@@
Limitation. The lightweight finetuning has limitations: 1) The 8k training examples represented a limited view of the behavior of our large user base. 2) For the cluster combinations that are hard to reason, LLM has low prediction confidence, indicated by the novel interest predictions that don’t have a logical connection to the users’ existing interests. This hurts the relevancy of the recommendation and, consequently, user satisfaction.
Figure 1: Hierarchical planning paradigm: the novelty LLM performs high-level planning for novel interest transitions, which are used to restrict the predictions of classic recommender models, and user feedback on these novel recommendations is aggregated to train a separate alignment LLM.
@@
To improve the relevancy of the novel interest prediction, we initially tried to increase the number of training examples. However, novelty metrics in A/B testing didn’t show sensitivity to this change. Furthermore, due to the LLM’s tendency to repeat training data (our analysis showed a 40% chance of repetition during inference), scaling to more training examples mined from the user history risked reinforcing the system feedback loop, impairing LLM’s ability to make novel recommendations.
@@
To align the LLM with user preference without amplifying the system feedback loop, we leverage live-traffic users’ feedback to LLM’s own recommendations, such as clicks, dwell time and repeated interaction, which is independent of the system behavior. We first tried the classic RLHF setup: RL fine-tune the novelty LLM directly with a reward model trained with user preference. However, this always resulted in the model quickly collapsing: 1) loss of controlled generation: after 5k steps, the LLM’s chance of predicting in the correct format drops from 99+to 2%; 2) Reward hacking: the model learned the high reward words, e.g., ’cat’, ’BTS’, ’toys’, etc, and frequently predicts those words. While RLHF is effective for free form text generation in conversation settings, it proved insufficient in structured tasks with strict format and content vocab requirements – the reward model cannot capture the nuanced task requirements and guide the RL finetuning process accordingly.
@@ 4 Method
To address the challenges in classic RLHF, we introduce an inference-time scaling method Brown et al. (2024) with a decoupled dual-specialization modeling approach. Instead of directly fine-tuning the policy model (i.e., the novelty model for planning the next cluster) through SFT or RLHF, we first performs independent sampling from the novelty model. This generates a diverse set of candidate interest clusters. Subsequently, the best-of-n clusters are selected using a separate alignment model trained on collective user feedback based on their likelihood to resonate with users.
@@ 4.1 Preference Alignment on User Feedback
Aggregating Collective Human Feedback. Through per-query logging inside our LLM-powered recommender serving live traffic (detailed in Section 3, ‘Novelty model’ hereafter), we collect users’ preferences on LLM’s predictions. Specifically, for each predicted cluster $C_{n}$, we log the cluster sequence {$C_{1}$, …, $C_{K}$} used to represent the user, and the user’s feedback on $C_{n}$ (e.g., positive playback, like, share, skip, etc). We then aggregate the feedback for each ({$C_{1}$, …, $C_{K}$}, $C_{n}$) pair, resulting in user preference training examples denoted as ({$C_{1}$, …, $C_{K}$}, $C_{n}$ , $L_{(1,k),n}$). Here, $L_{(1,...,k),n}$ represents the aggregated user feedback score (e.g. like rate, share rate) for this particular interest cluster transition – that is, serving interest cluster $C_{n}$ to a user with historical viewing pattern represented by {$C_{1}$, …, $C_{K}$}).
We then post-process the aggregated feedback to: 1) normalize the feedback score, which can be skewed towards very small values because the feedback signals, e.g. like, share, etc, are sparse. 2) filter cluster transition pairs with little user feedback. 3) round the feedback score to a fixed interval to account for margin of error in the aggregated stats.
@@
Alignment Reward Model Training. To align with collective user feedback, we trained an "alignment model"(a reward model) to score the users’ affinity to $C_{n}$ given their watch history. The alignment model is training using a cross-entropy loss between its prediction and the user’s actual aggregated engagement metric (i.e., positive playback rate). This alignment model is an LLM with the last layer being a linear projection layer.
@@
In Figure 2, we showcase a sample prompt describing users with {photography, Visual Effects, Special effects} and {dogs} interest clusters(assuming $K=2$). Collectively, those users prefer label 1({wildlife, nature documentary}) over label 2 ({boats}) as expressed in the feedback scores. Given two new labels, the trained alignment model also effectively assigns high preference score to {cats, video game, internet meme} over less relevant next interest cluster. These intuitive examples demonstrate the feasibility and potential of the alignment training.